OpenStudy (anonymous):

show whether the integral converges or diverges is the integral finite? \[\int_1^\infty \frac{1}{2x - 1} \mathrm{d}x\]

OpenStudy (anonymous):

diverges for sure

OpenStudy (anonymous):

in order for it to converge, the degree of the denominator must be larger than the degree of the numerator by more than one

OpenStudy (anonymous):

if you integrate that thing, you will get the log of something when you take the limit as \(x\to \infty\) of the log of whatever, it will go to infinity

OpenStudy (anonymous):

\[log_{something} ( \inf ) = \inf\]

OpenStudy (anonymous):

for example \[\int_1^{\infty}\frac{x+1}{x^2-2}dx\] diverges, whereas \[\int_1^{\infty}\frac{x+1}{x^3-2}dx\] converges

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

\[anybase^{\inf} = \inf\]is what the logarithm x->inf is about?

OpenStudy (anonymous):

^ "when you take the limit as x→∞ of the log of whatever, it will go to infinity"

OpenStudy (anonymous):

\[\lim_{x \rightarrow \inf} log_{whatever} ( x) = \inf\]is the same statement as \[whatever^{\inf} = x,x\rightarrow \inf\]?